OpenStudy (anonymous):

Solve the differential equation y dx+(2xy-e^(-2y))dy=0. I know the integrating factor is e^2y/y but what do I do next?

OpenStudy (zale101):

\[\frac{ e^{2y} }{ y }y dx+\frac{ e^{2y} }{ y }(2xy-e^{-2y})dy=0\] scale equation and simplify

OpenStudy (zale101):

can you do that?

OpenStudy (anonymous):

Wait a minute.

OpenStudy (zale101):

ok :)

OpenStudy (anonymous):

I got e^2y dx+2x(e^2y)-1/y dy=0, now what?

OpenStudy (anonymous):

Do I integrate now?

OpenStudy (zale101):

oh i'm sorry for the late reply

OpenStudy (zale101):

:\[M_1(x,y)=e ^{2y} and N(x,y)=2xe ^{2y}-1/y\]

OpenStudy (anonymous):

I already know that. From e^2y dx+2x(e^2y)-1/y dy=0, do I need to integrate? The answer is xe^2y-ln abs(y)=c, y=0. How do I get that?

OpenStudy (zale101):

\[M_1(x,y)=\int\limits_{}^{} M (x,y) dx=\int\limits_{}^{}e^{2y} dx= xe ^{2y}\]

OpenStudy (anonymous):

I don't get what you wrote.

OpenStudy (zale101):

i just confirmed

OpenStudy (zale101):

but looks like,constant of integration can't be added on what i wrote above

OpenStudy (anonymous):

Why is y=0 part of the answer?

OpenStudy (zale101):

this is not the answer?

OpenStudy (anonymous):

The answer is xe^2y-ln abs(y)=c, y=0. But I don't know how to get y=0 as the answer.

OpenStudy (zale101):

X is not dependent

OpenStudy (anonymous):

How so?

OpenStudy (zale101):

\[M_1(x,y) + \int\limits_{}^{}[ N(x,y)-\frac{ ∂ }{ ∂y } M_1(x,y)] dy=xe^{2y}-\int\limits_{}^{}\frac{ 1 }{ y }dy=xe ^{2y}-In|y|+C=0\]

OpenStudy (zale101):

btw, C is an arbitrary constant

OpenStudy (zale101):

what, it didn't show fully

OpenStudy (zale101):

\[\frac{ 1 }{ y }dy=xe ^{2y}-In|y|+C=0\]

OpenStudy (anonymous):

I still don't know how to get y=0.

OpenStudy (zale101):

Basically, the original equation in the form dy/dx = f(x, y)

OpenStudy (zale101):

e^integral(2dy) = e^(2y) (x*e^(2y))' = e^(2y)*e^(-2y)/y = 1/y Integrate:x*e^(2y) = ln(y) + c then you'll get for the solution: x(y) = (ln(y) + c) * e^(-2y)

OpenStudy (zale101):

M=e^2y and N=2xe^2y-1/y, that means My=Nx=2e^2y